OpenStudy (anonymous):

Find 3rd degree function with real coefficients, if the roots are -3, 1-2i

OpenStudy (campbell_st):

for the roots you'll have (x - 1)^2 = -4 to check \[x -1 = \sqrt{-4}\] then \[x - 1 = \sqrt{2i^2}\] \[x - 1 = \pm 2i\] so \[x = 1 \pm 2i\] so going back (x -1)^2 = 4 is the same as x^2 - 2x + 1 - 4 so (x^2 - 2x - 3) is a quadratic factor the other factor is (x + 3) so the factored polynomial is \[P(x) = (x + 3)(x^2 - 2x - 3) \] just distribute to get the cubic equation